# Impact of Annealing and Nanostructuring on Properties of NV Centers Created by Different Techniques

### v1

@@ Abstract
Nitrogen-vacancy (NV) centers in diamonds have been an epicenter of research for diverse applications in quantum technologies. It is therefore imperative that their fabrication techniques are well understood and characterized for the technological scalability of these applications. A comparative study of the optical and spin properties of NVs created by ion implantation and chemical vapor deposition delta-doping is thus presented, combined with an investigation on the impact of annealing in vacuum at different temperatures. In addition, nanopillars are fabricated with electron beam lithography and reactive ion etching for enhanced photon collection efficiency. An extensive combination of characterization techniques is employed. Notably, the smallest nanopillars present fluorescence enhancements of factor around 50, compared to the unstructured regions. Annealing is also demonstrated to increase the optical contrast between the NVs’ electronic states, the coherence and relaxation times both in bulk as in pillars. Regarding the NV preparation technique, the delta-doping is shown to create NVs with less lattice defects and strain compared to implantation.
@@ I Introduction
The NV center is a point defect in the diamond lattice composed of a nitrogen atom replacing a carbon atom adjacent to a vacancy, featuring unique optical and spin properties. A groundbreaking demonstration in the early 2000s showed that the electron spin of NV centers could be optically manipulated and read out at room temperature F.Jelezko2 , marking the beginning of their implementation in quantum sensing sensing_NV1 ; sensing_NV2 ; sensing_NV3 , communications communication_NV1 and computing F.Jelezko ; computing_NV1 ; computing_NV2 . Subsequent research demonstrated the ability of NV centers to measure weak magnetic fields at the nanoscale, highlighting their potential in quantum magnetometry and nanoscale sensing Balasubramanian ; sensor . All these findings motivate further research on the NV properties and the precise characterization of its preparation techniques, in order to reach technological scalability for these applications.
@@
Depending on the application, different techniques are preferred for creation of NV centers. For instance, nitrogen implantation is favored for creating single NV centers used in quantum computing, since it offers precise control over their depth and positioning ArifulHaque ; FedericoGorrini . Chemical vapor deposition (CVD) is commonly preferred for creating NV-rich diamonds for sensing and imaging applications TLuo . Furthermore, heavily doped nitrogen delta layers are created by microwave plasma-assisted CVD with a rapid gas switching and a laminar gas flow to create sharp interfaces between doped and undoped diamond layers with thicknesses down to 2-3 nm deltaNVVikharev ; deltaNVBogdanov . Each of these methods has advantages and disadvantages. In the case of ion implantation, a precise control over the concentration and localization of the implanted ions can be achieved varying the energy and dose of implanted ions, which affect the NV properties. The implantation process, however, causes significant damage of the diamond lattice, deteriorating the relaxation times of the NV centers Jelezko3 . Contrastingly in CVD NV-diamonds, the NV centers have better coherence properties, in the expense of a lower control over their localization and concentration AcostaHemmer .
@@
Another important point for applications in quantum technology is the incorporation of NV centers in photonic structures for enhancement of improving photon collection efficiency nanowire , which is one of the requirements for a high-fidelity NV state readout DominikIrber . Finally, annealing also plays a crucial role for the NV centers creation and their properties annealingLPHT ; annealing1 ; annealing2 not only by improving the conversion efficiency of P1 centers to NVs, but also changing the local charge distribution and the diamond lattice strain around the NVs by enabling diffusion of charged defects charge_diffusion and relieving the lattice strain.
@@ II.1 CVD $\delta$-doping
Nitridation process with pure N2 plasma in a microwave plasma (MWP) CVD setup was applied, followed by a short in situ diamond CVD overgrowth to create a nitrogen $\delta$-doped layer technion ; technion2 . Electronic grade diamond 4.0 $\times$ 4.0 $\times$ 0.5 mm (Element Six) was used as a substrate. It was initially exposed to MWP-H cleaning of 30 min at 6 kW plasma power and 100 Torr (133.32 mbar) working pressure. Then the plasma was changed to pure N2 (99.999%) and treated for 30 min at 10 Torr (13.33 mbar) and 1 kW. After that, the chamber was evacuated and diamond overgrowth was performed for 10 min at 100 Torr (133.32 mbar), 6 kW, 2% CH4 in H2, and 950C substrate temperature. The thickness of the overgrown layer was about 800 nm considering the typical growth rate at the above mentioned conditions. Previous studies have shown the formation of N delta layer in diamond as revealed by sharp SIMS profiles of CN- in the spectra technion ; technion2 .
@@ II.4 Annealing Under Vacuum Condition
Previous studies have shown plasma-induced lattice damages and stress after ICP-RIE with O2 plasma for diamond structuring, resulting in reduction of the photoluminescence emission and degradation of the spin properties of NV centers located in the vicinity of the etched surface ICPdamageNVnearSurface ; ICPdamageNVnearSurface2 ; ICPdamageNVnearSurface3 . Annealing of the samples can improve the properties of NV centers, as demonstrated by Meng et al. annealingLPHT . The diamond samples were annealed at pressures below $5\times 10^{-6}$ mbar after nanostructuring in a home-made furnace. Details of the annealing protocols can be found in the Supplementary Material (Sec. S2).
@@ II.5 Characterization of Diamond Samples
The properties of the NV centers were characterized in a custom-made confocal microscope, as similarly described in setup_1 ; setup_2 . The NV ensembles were optically excited with a scanning green laser of 518 nm wavelength focused on the diamond through an air objective, also used to collect the red fluorescence to determine the $m_{s}$ state. This resulted in a laser spot diameter of roughly 660 nm nanopillars_kseniia . The fluorescence was then filtered, detected by an avalanche photo diode and counted. Alternatively, in photoluminescence experiments, the fluorescence was detected by a spectrometer. In saturation curve measurements, the laser power was varied and measured with a powermeter sensor, while in the other experiments the laser power was kept at 0.5 mW for efficient polarization of the $m_{s}$ state.
Microwave fields were generated either by a continuous source or by an arbitrary waveform generator, depending on the experiment. In all cases, the microwave was amplified and transmitted to the NVs through a thin copper wire of 20 $\mu$m diameter. For the laser spot size under consideration, the resulting microwave field $\vec{B}_{1}(t)$ was homogeneous in good approximation. A static magnetic field $\vec{B}_{0}$ was applied by a permanent magnet resulting in a small intensity of about $|\vec{B}_{0}|\approx 3.6$ mT, which was also homogeneous within the laser spot size. Finally, Qudi software qudi was used for hardware control and data acquisition.
@@ III.2 Photoluminescence
Photoluminescence measurements were performed in the bulk region non structured area of all three diamond samples before annealing. Fig. 3 (a) shows the spectrum from the $\delta$-doped sample and (b) from the (111)-implanted one. The (100)-implanted sample exhibits a spectrum equivalent to the (111) and thus is not shown here. Primarily, all samples have the characteristic zero-phonon line (ZPL) from NV- at 637.2 nm, indicating the effective creation of NV centers. In the implanted samples, the ZPL from NV0 is also observed at 575 nm. On the other hand, the $\delta$-doped sample presents a strong Raman emission from the diamond host below 550 nm raman_diamond , which saturates the counts at higher wavelengths, thus, requiring the use of 600 nm long pass filter blocking the NV0 ZPL. Finally, an intense emission from silicon-vacancy (SiV) centers is also observed at 737 nm in the spectrum of the $\delta$-sample, which could be related to their creation from Si impurities during the CVD overgrowth of the sample.
@@ III.3 Fluorescence Mapping and Saturation Curves
The successful fabrication of the nanopillars is further confirmed by the fluorescence mapping as shown in Fig. 4, for the $\delta$-doped sample before annealing. On the larger $200\times 200$ $\mu$m2 scan [Fig. 4 (a)], letter and number markers are visible, representing areas (A-E) with different nominal pillar diameters given in Table S1. The shadow cast by the microwave wire antenna is also seen across $x$ at $y\approx 100$ $\mu$m. On Fig. 4 (b), a smaller scale scan is presented over the B marker pillars, resulting in increased fluorescence. In addition, a small number of defects in the fabrication is also observed. The other two samples show analogous fluorescence mapping before annealing.
@@
Figure 4: Fluorescence mapping of $\delta$-doped sample before annealing for (a) large scan area and (b) zoomed in the pillars from marker B area. In the larger range, it is possible to visualize the markers indicating the regions with pillars of different diameters, whereas in the smaller region, the high fluorescence arising from the pillars can be observed.
@@
Figure 5: Saturation curves of the bulk region after annealing at 1400C in the (a) $\delta$-doped, (b) (100)-implanted and (c) (111)-implanted samples. The (111)-implanted sample has a fluorescence of about four times smaller than the other two, given by a less efficient NV fabrication yield. From the saturation curves, the enhancement factors of the pillars were calculated for the (d) $\delta$-doped, (e) (100)-implanted and (f) (111)-implanted samples. The $\delta$-doped and (111)-implanted samples show significant increase in the area normalized fluorescence of the pillars, which decreases with the pillar diameter, whereas the pillars in the (100)-implanted one have fluorescence comparable with the unstructured region.
@@
Photon collection efficiency properties of the photonic nanostructures can be obtained from fluorescence saturation curve measurements, where the fluorescence counts are measured as a function of the laser power nanopillars_kseniia ; sat_function . This way, saturation curves were obtained in the pillars and in the bulk regions, as shown in Fig. 5 (a), (b) and (c) for the bulk region after annealing at 1400C in the $\delta$-doped, (100)-implanted and (111)-implanted samples, respectively. The $\delta$-doped and (100)-implanted samples exhibit comparable fluorescence in the bulk, whereas the (111)-implanted one has fluorescence about four times smaller, which could be related to a less efficient NV yield by the implantation at this crystallographic orientation nanopillars_kseniia .By the $\delta$-doped sample the saturation curves were measured before and after each annealing step. By the implanted samples a bleaching effect was observed before annealing, marked by a decay of the fluorescence over time, thus making these measurements unfeasible.
@@
To calculate the enhancement factor from the pillars, the fluorescence was fitted as a function of the laser power according to the definition of P. Siyushev et al. sat_function
@@
where $P$ is the laser power, $\mathcal{F}_{\infty}$ the fluorescence at saturation, $P_{sat}$ the saturation power and $C$ a constant offset from the background. The enhancement factor is given by $\mathcal{F}_{\infty}$ parameter divided by the reference value of the unstructured bulk for the same annealing condition in each sample, then normalized by the pillar average area measured by SEM (Sec. III.1) or by the area of the laser spot (with an estimated diameter of 660 nm) in the case of the bulk region nanopillars_kseniia . Additionally, the uncertainty of the enhancement factor is calculated from error propagation of the fit uncertainties and experimental deviation of the pillars diameters.
The $\delta$-doped and (111)-implanted samples show significant increase (slightly higher by the first one) in the area normalized fluorescence of the pillars, reaching enhancement values around 50 for the smallest pillars. As the pillars diameter increases the enhancement factors decrease, approaching the value of the unstructured bulk region. Conversely, in the (100)-implanted diamond the fluorescence in the smallest pillars is slightly below that of the bulk, and increases with the pillar diameter. This can be attribute to some graphitization of the sample caused by annealing, affecting the NV optical properties. Lastly, the annealing of the $\delta$-doped sample also induces a small decline in the enhancement factors of the smallest pillars.
@@ III.4 Optically Detected Magnetic Resonance
with free parameter of $a_{i}$ as the intensity of each peak, $\delta\omega_{i}$ the width, $\omega_{i}$ the position and $C$ the total offset, typically 1. In turn, the total contrast of each spectrum as shown in Fig. 6 (a) is defined by the average intensity $\sum_{i}a_{i}/6$, whereas the uncertainty is calculated from the covariance of the fit parameters. Notably, the contrast increases significantly with the annealing by a factor of almost three in the $\delta$-doped and (111)-implanted samples, and of 5.3 in the (100)-implanted one. The intermediate annealing step at 1200C of the $\delta$-doped sample also confirms this trend. In regards to the NV creation technique, the $\delta$-doped sample showed lower contrast enhancement after annealing than the other two.

### v2

@@ Abstract
Nitrogen-vacancy (NV) centers in diamonds have been an epicenter of research for diverse applications in quantum technologies. It is therefore imperative that their fabrication techniques are well understood and characterized for the technological scalability of these applications. A comparative study of the optical and spin properties of NVs created by ion implantation and chemical vapor deposition delta-doping is thus presented, combined with an investigation on the impact of annealing in vacuum at different temperatures. In addition, nanopillars are fabricated by electron beam lithography and reactive ion etching for enhanced photon collection efficiency. An extensive combination of characterization techniques is employed. Notably, the smallest nanopillars present fluorescence enhancements of factor around 50, compared to the unstructured regions. Annealing is also demonstrated to increase the optical contrast between the NVs’ electronic states, the coherence and relaxation times both in bulk as in pillars. Regarding the NV preparation technique, the delta-doping is shown to create NVs with less lattice defects and strain compared to implantation.
@@ I Introduction
The NV center is a point defect in the diamond lattice composed of a nitrogen atom replacing a carbon atom adjacent to a vacancy, featuring unique optical and spin properties. A groundbreaking demonstration in the early 2000s showed that the electron spin of NV centers could be optically manipulated and read out at room temperature Jelezko et al. (2004a), marking the beginning of their implementation in quantum sensing Chen et al. (2022); Aslam et al. (2023); Fujiwara et al. (2020), communications Ruf et al. (2021) and computing Jelezko et al. (2004b); Waldherr et al. (2014); Stas et al. (2022). Subsequent research demonstrated the ability of NV centers to measure weak magnetic fields at the nanoscale, highlighting their potential in quantum magnetometry and nanoscale sensing Balasubramanian et al. (2008); Taylor et al. (2008). All these findings motivate further research on the NV properties and the precise characterization of its preparation techniques, in order to reach technological scalability for these applications.
@@
Depending on the application, different techniques are preferred for creation of NV centers. For instance, nitrogen implantation is favored for creating single NV centers used in quantum computing, since it offers precise control over their depth and positioning Haque and Sumaiya (2017); Gorrini and Bifone (2023). Chemical vapor deposition (CVD) is commonly preferred for creating NV-rich diamonds for sensing and imaging applications Luo et al. (2022). Furthermore, heavily doped nitrogen delta layers are created by microwave plasma-assisted CVD with a rapid gas switching and a laminar gas flow to create sharp interfaces between doped and undoped diamond layers with thicknesses down to 2-3 nm Vikharev et al. (2016); Bogdanov et al. (2020). Each of these methods has advantages and disadvantages. In the case of ion implantation, a precise control over the concentration and localization of the implanted ions can be achieved varying the energy and dose of implanted ions, which affect the NV properties. The implantation process, however, causes significant damage of the diamond lattice, deteriorating the relaxation times of the NV centers Pezzagna et al. (2010a). One possible solution to minimize the lattice damage would be to keep the sample at high temperature during implantation. Thus, the defects created by the ions could diffuse and annihilate when reaching the diamond surface. Contrastingly in CVD NV-diamonds, the NV centers have better coherence properties, in the expense of a lower control over their localization and concentration Acosta and Hemmer (2013). In addition to these two mostly used techniques, NVs can be also created by direct laser writing as recently shown Addhya et al. (2024).
@@
Another important point for applications in quantum technology is the incorporation of NV centers in photonic structures for enhancement of improving photon collection efficiency Hausmann et al. (2010), which is one of the requirements for a high-fidelity NV state readout Irber et al. (2021). Finally, annealing also plays a crucial role for the NV centers creation and their properties fei Meng et al. (2008); Naydenov et al. (2010); Pezzagna et al. (2010b) not only by improving the conversion efficiency of P1 centers to NVs, but also changing the local charge distribution and the diamond lattice strain around the NVs by enabling diffusion of charged defects Lühmann et al. (2019) and relieving the lattice strain.
@@ II.1 CVD $\delta$-doping
Nitridation process with pure N2 plasma in a microwave plasma (MWP) CVD setup was applied, followed by a short in situ diamond CVD overgrowth to create a nitrogen $\delta$-doped layer Jaffe et al. (2020); Kuntumalla et al. (2021). Electronic grade diamond 4.0 $\times$ 4.0 $\times$ 0.5 mm (Element Six) was used as a substrate. It was initially exposed to MWP-H cleaning of 30 min at 6 kW plasma power and 100 Torr (133.32 mbar) working pressure. Then the plasma was changed to pure N2 (99.999%) and treated for 30 min at 10 Torr (13.33 mbar) and 1 kW. After that, the chamber was evacuated and diamond overgrowth was performed for 10 min at 100 Torr (133.32 mbar), 6 kW, 2% CH4 in H2, and 950C substrate temperature. The thickness of the overgrown layer was about 800 nm considering the typical growth rate at the above mentioned conditions. Previous studies have shown the formation of N delta layer in diamond as revealed by sharp SIMS profiles of CN- in the spectra Jaffe et al. (2020); Kuntumalla et al. (2021).
@@ II.4 Annealing Under Vacuum Condition
Previous studies have shown plasma-induced lattice damages and stress after ICP-RIE with O2 plasma for diamond structuring, resulting in reduction of the photoluminescence emission and degradation of the spin properties of NV centers located in the vicinity of the etched surface Cui et al. (2015); Kim et al. (2014); Fávaro de Oliveira et al. (2015). Annealing of the samples can improve the properties of NV centers, as demonstrated by Meng et al. fei Meng et al. (2008). The diamond samples were individually annealed at pressures below $5\times 10^{-6}$ mbar after nanostructuring in a home-made furnace. Details of the annealing protocols can be found in Sec. S2.
@@ II.5 Characterization of Diamond Samples
The properties of the NV centers were characterized in a custom-made confocal microscope, as similarly described in Tsunaki et al. (2024a); Panadero et al. (2024). The NV ensembles were optically excited with a scanning green laser of 518 nm wavelength focused on the diamond through an air objective, also used to collect the red fluorescence to determine the $m_{s}$ state. This resulted in a laser spot diameter of roughly 660 nm Volkova et al. (2022). The fluorescence was then filtered, detected by an avalanche photo diode and counted. Alternatively, in photoluminescence experiments, the fluorescence was detected by a spectrometer. In saturation curve measurements, the laser power was varied and measured with a powermeter sensor, while in the other experiments the laser power was kept at 0.5 mW for efficient polarization of the $m_{s}$ state.
Microwave fields were generated either by a continuous source or by an arbitrary waveform generator, depending on the experiment. In all cases, the microwave was amplified and transmitted to the NVs through a thin copper wire of 20 $\mu$m diameter. For the laser spot size under consideration, the resulting microwave field $\vec{B}_{1}(t)$ was homogeneous in good approximation. A static magnetic field $\vec{B}_{0}$ was applied by a permanent magnet resulting in a small intensity of about $|\vec{B}_{0}|\approx 3.6$ mT, which was also homogeneous within the laser spot size. Finally, Qudi software Binder et al. (2017) was used for hardware control and data acquisition.
@@ III.2 Photoluminescence
Photoluminescence measurements were performed in the bulk region non structured area of all three diamond samples before annealing. Fig. 3 (a) shows the spectrum from the $\delta$-doped sample and (b) from the (111)-implanted one. The (100)-implanted sample exhibits a spectrum equivalent to the (111) and thus is not shown here. Primarily, all samples have the characteristic zero-phonon line (ZPL) from NV- at 637.2 nm, indicating the effective creation of NV centers. In the implanted samples, the ZPL from NV0 is also observed at 575 nm. On the other hand, the $\delta$-doped sample presents a strong Raman emission from the diamond host below 550 nm Petkov et al. (2013), which saturates the counts at higher wavelengths, thus, requiring the use of 600 nm long pass filter blocking the NV0 ZPL. Finally, an intense emission from silicon-vacancy (SiV) centers is also observed at 737 nm in the spectrum of the $\delta$-sample, which could be related to their creation from Si impurities during the CVD overgrowth of the sample.
@@ III.3 Fluorescence Mapping and Saturation Curves
The successful fabrication of the nanopillars with incorporated NV centers is further confirmed by the fluorescence mapping as shown in Fig. 4 for the $\delta$-doped sample before annealing. Fig. 4 (a) and (b) show $200\times 200$ $\mu$m2 scans, where letter and number markers are visible, representing areas E to C and C to A, respectively, with different nominal pillar diameters given in Tab. S1. The shadow cast by the microwave wire antenna is also seen across $x$ at $y\approx 100$ $\mu$m. Due to the large scanned area and the resulting tilt of the diamond surface, not all regions can be optically focused simultaneously and some regions appear brighter than others. On Fig. 4 (c), a smaller scale scan is presented over the B marker pillars, resulting in increased fluorescence. In addition, a small number of defects in the fabrication is also observed, as well as some missing pillars of the smallest diameter in region E. The other two samples show analogous fluorescence mapping before annealing.
@@
Figure 4: Fluorescence mapping of $\delta$-doped sample before annealing for (a) large scan areas showing regions E to C with small pillars, (b) regions C to A with larger pillars and (c) zoomed in the pillars from marker B area.
@@
Figure 5: Area normalized saturation curves of the bulk region after annealing at 1400C in the (a) $\delta$-doped, (b) (100)-implanted and (c) (111)-implanted samples. The (111)-implanted sample has a fluorescence of about four times smaller than the other two, given by a less efficient NV fabrication yield. From the saturation curves, the enhancement factors of the pillars were calculated for the (d) $\delta$-doped, (e) (100)-implanted and (f) (111)-implanted samples. The $\delta$-doped and (111)-implanted samples show significant increase in the area normalized fluorescence of the pillars, which decreases with the pillar diameter, whereas the pillars in the (100)-implanted one have fluorescence comparable with the unstructured region.
@@
Even though the fluorescence mapping gives qualitative information about the photon collection efficiency properties of the photonic nanostructures, a more precise quantitative comparison could be obtained from fluorescence saturation curve measurements. In this case, the fluorescence counts are measured as a function of the laser power Volkova et al. (2022); Siyushev et al. (2010) and then normalized by the area. The latter is determined by the laser spot size for the bulk (with an estimated diameter of 660 nm Volkova et al. (2022)) and by the diameters of the pillars measured by SEM. This way, saturation curves were obtained in the pillars and in the bulk regions, as shown in Fig. 5 (a), (b) and (c) for the bulk region after annealing at 1400C in the $\delta$-doped, (100)-implanted and (111)-implanted samples, respectively. The $\delta$-doped and (100)-implanted samples exhibit comparable fluorescence in the bulk, whereas the (111)-implanted one has fluorescence about four times smaller, which could be related to a less efficient NV yield by the implantation at this crystallographic orientation Volkova et al. (2022).By the $\delta$-doped sample the saturation curves were measured before and after each annealing step. By the implanted samples a bleaching effect was observed before annealing, marked by a decay of the fluorescence over time, thus making these measurements unfeasible. Further saturation curves for the pillar regions of the three studied samples are shown in Figs. S3 and S4 in the Supplementary Material.
@@
To calculate the enhancement factor from the pillars, the area normalized fluorescence was fitted as a function of the laser power according to the definition of P. Siyushev et al. Siyushev et al. (2010)
@@
where $P$ is the laser power, $\mathcal{F}_{\infty}$ the fluorescence at saturation, $P_{sat}$ the saturation power and $C$ a constant offset from the background. The enhancement factor is given by $\mathcal{F}_{\infty}$ parameter divided by the reference value of the unstructured bulk for the same annealing condition in each sample, then normalized by the pillar average area measured by SEM (Sec. III.1) or by the area of the laser spot (with an estimated diameter of 660 nm) in the case of the bulk region Volkova et al. (2022). Additionally, the uncertainty of the enhancement factor is calculated from error propagation of the fit uncertainties and experimental deviation of the pillars diameters.
The $\delta$-doped and (111)-implanted samples show significant increase (slightly higher by the first one) in the area normalized fluorescence of the pillars, reaching enhancement values around 50 for the smallest pillars, although they exhibit less intense absolute fluorescence than the larger ones as seen in Fig. 4. As the pillars diameter increases the enhancement factors decrease, approaching the value of the unstructured bulk region. Conversely, in the (100)-implanted diamond the fluorescence in the smallest pillars is slightly below that of the bulk, and increases with the pillar diameter. This can be attribute to some graphitization of the sample caused by annealing, affecting the NV optical properties. The graphitization could be due to surface imperfections which vary for the different samples. Lastly, the annealing of the $\delta$-doped sample also induces a small decline in the enhancement factors of the smallest pillars. The inhomogeneity in the fluorescence enhancement factors between the different samples and diameters could be ascribed not only to surface defects leading to partial graphitization, but also to a great extent to the crystal quality. The overgrown nitrogen $\delta$-doped sample shows much stronger Raman signal (Sec. III.2) related to diamond, compared to the other two samples, suggesting a crystal lattice with less defects. After the annealing some of the defects in the implanted samples are removed, leading to an improved crystal quality of the pillars.
@@ III.4 Optically Detected Magnetic Resonance
with free parameter of $a_{i}$ as the intensity of each peak, $\delta\omega_{i}$ the width, $\omega_{i}$ the position and $C$ the total offset, typically 1. In turn, the total contrast of each spectrum as shown in Fig. 6 (a) is defined by the average intensity $\sum_{i}a_{i}/6$, whereas the uncertainty is calculated from the covariance of the fit parameters. Notably, the contrast increases significantly with the annealing by a factor of almost three in the $\delta$-doped and (111)-implanted samples, and of 5.3 in the (100)-implanted one. The intermediate annealing step at 1200C of the $\delta$-doped sample also confirms this trend. In regards to the NV creation technique, the $\delta$-doped sample showed lower contrast enhancement after annealing than the implanted samples. This could be related to removal of some fluorescent defects and/or bettered stabilized NV charge state after the annealing Naydenov et al. (2010), better exhibited by implanted samples.